# Beyond Detection: Exploring Evidence-based Multi-Agent Debate for Misinformation Intervention and Persuasion

### v1

@@ Abstract
Multi-agent debate (MAD) frameworks have emerged as promising approaches for misinformation detection by simulating adversarial reasoning. While prior work has focused on detection accuracy, it overlooks the importance of helping users understand the reasoning behind factual judgments and develop future resilience. The debate transcripts generated during MAD offer a rich but underutilized resource for transparent reasoning. In this study, we introduce ED2D, an evidence-based MAD framework that extends previous approach by incorporating factual evidence retrieval. More importantly, ED2D is designed not only as a detection framework but also as a persuasive multi-agent system aimed at correcting user beliefs and discouraging misinformation sharing. We compare the persuasive effects of ED2D-generated debunking transcripts with those authored by human experts. Results demonstrate that ED2D outperforms existing baselines across three misinformation detection benchmarks. When ED2D generates correct predictions, its debunking transcripts exhibit persuasive effects comparable to those of human experts; However, when ED2D misclassifies, its accompanying explanations may inadvertently reinforce users’ misconceptions, even when presented alongside accurate human explanations. Our findings highlight both the promise and the potential risks of deploying MAD systems for misinformation intervention. We further develop a public community website to help users explore ED2D, fostering transparency, critical thinking, and collaborative fact-checking.
@@ Introduction
Misinformation presents a persistent threat to online discourse, public trust, and democratic institutions (10.1145/3733567.3735568; 10.1145/3733567.3735566). In response, researchers have developed a range of computational approaches to automatically detect misleading content. Among these, multi-agent debate (MAD) frameworks have recently gained attention for their ability to simulate adversarial reasoning (liang-etal-2024-encouraging; li-etal-2024-improving-multi), where large language model (LLM) agents engage in structured argumentation to expose factual inconsistencies. These multi-agent systems leverage the complementary strengths of competing perspectives and often yield more robust judgments than a single classifier.
While prior studies employ MAD frameworks for post-verification tasks that focus on identifying the falsehood of claims, they often neglect the more fundamental challenge of proactive misinformation intervention. In the real world, users are active reasoners rather than passive recipients, requiring persuasive explanations and the capacity to resist future misinformation. (10.1145/3706598.3713408). Simply labeling a claim as false is insufficient to foster resilience against misinformation (LYU2025130536). Drawing on the adage that the Truth Becomes Clearer Through Debate (han2025d2d; 10.1145/3726302.3730092), we argue that effective misinformation interventions should prioritize transparent reasoning, evidence grounding, and persuasive debunking.
In this study, we introduce ED2D, an evidence-based MAD framework that extends the Debate-to-Detect (D2D) framework (han2025d2d) by incorporating an evidence retrieval module. It automatically identifies key entities and relations within a claim, retrieves factual information from external sources, and assesses the stance of the retrieved evidence with respect to the claim. By integrating verifiable facts into the argumentative process, ED2D mitigates hallucinations and achieves superior performance on three standard misinformation detection benchmarks.
We further evaluate ED2D’s persuasive effect on human beliefs using Snopes25, a real-world benchmark consisting of fact-checks authored by professional editors from January to June 2025. Snopes25 is collected after the training cut-off of GPT-4o, thereby reducing the risk of data leakage and better reflecting current misinformation trends. Using this dataset, we compare the effects of ED2D-generated debunks and human expert reports on belief correction and sharing intention. We also analyze failure cases in which ED2D misjudges claims yet still successfully persuades users, highlighting potential risks associated with the deployment of MAD systems.
In summary, this study makes the following contributions:
@@
Early approaches focused on fine-tuning pretrained transformers such as BERT for binary veracity classification tasks (10.1007/978-3-030-73696-5_11). While these approaches perform well on benchmark datasets, they often lack interpretability and generalize poorly to novel claims (pelrine-etal-2023-towards). To overcome these limitations, recent work has shifted toward prompting-based paradigms using LLMs. Zero-shot and few-shot prompting enable flexible adaptation to misinformation detection tasks without the need for explicit fine-tuning (10.1145/3583780.3615015). Moreover, techniques such as Chain-of-Thought (Weicot) and Self-Reflection (madaan2023selfrefine) enhance reasoning transparency by enabling models to decompose complex judgments into intermediate steps or engage in iterative self-critique. However, as these approaches depend entirely on internal model knowledge, they are prone to hallucinations and factual errors (10.1145/3703155). To address these reliability concerns, an emerging line of research incorporates external evidence into the reasoning process. For example, Retrieval-Augmented Generation (RAG) has been widely adopted to ground LLM outputs in verifiable content. Nevertheless, most existing RAG systems still rely on a fixed knowledge base, which limits their effectiveness in open-domain misinformation detection scenarios (wang2025evidence).
@@
MAD frameworks have gained traction as an approach to improve the deliberative quality of LLM outputs by simulating adversarial interactions between agents with opposing views (Du2024; liang-etal-2024-encouraging; li-etal-2024-improving-multi). Drawing inspiration from human deliberation and dialectical reasoning, such frameworks allow agents to challenge assumptions, correct errors, and elicit deeper justifications.
@@
Beyond accurate classification, effective misinformation mitigation requires persuasive interventions that influence user beliefs and behaviors (Ecker2022). Research in cognitive psychology has explored the design of corrective messages, demonstrating that informative and clear explanations can reduce belief in falsehoods and discourage the sharing intention (10.1145/3449092). Factors such as emotion, epistemic trust, and narrative coherence have been shown to moderate the effectiveness of fact-checking efforts (Pennycook2021; Scherer2021).
Professional fact-checking organizations, such as Snopes and PolitiFact, consistently produce structured and evidence-based rebuttals that represent the current gold standard in combating online misinformation. However, expert-written fact-checks are often costly and inherently lack scalability. Consequently, there is growing interest in automating persuasive debunking using LLMs (Salvi2025; schoenegger2025largelanguagemodelspersuasive). However, few studies have directly compared the persuasive efficacy of AI-generated versus expert-written explanations.
@@
Figure 2 presents the architecture of ED2D. The framework builds upon the five-stage debate structure: Opening Statement, Rebuttal, Free Debate, Closing Statement, and Judgment. At the core of the system is the Agent Layer, which consists of two debating teams—Affirmative and Negative—each composed of four agents. These agents are assigned domain-specific profiles relevant to the input claim and are fixed to either a ”True” or ”Fake” stance. Each team collaboratively constructs arguments to support or refute the claim’s veracity.
The debate is evaluated by a panel of judge agents, who observe the full dialogue and score it across five dimensions: Factuality, Source Reliability, Reasoning Quality, Clarity, and Ethical Considerations. Each dimension is assessed using a complementary scoring scheme, where paired scores sum to seven, ensuring that no tie can occur. The aggregated scores result in a definitive classification of the claim as either REAL or FAKE.
The Orchestrator Layer governs the overall debate process, assigning roles, scheduling turns across the five structured stages, and maintaining a compressed shared dialogue memory. To address the context length limitations of LLMs, it performs stage-wise context compression, distilling salient content into concise summaries that guide subsequent reasoning stages. This mechanism ensures continuity and coherence throughout the multi-turn debate.
Evidence Retrieval and Integration
  1. 1.
    Entity and Relation Extraction: Using in-context prompting with LLMs, the system identifies up to five salient entities, concepts, or relations from the input claim.
  2. 2.
  3. 3.
  4. 4.
Judgment and Output
In the final stage, ED2D generates two outputs: a binary veracity label and a structured summary of the debate. The summary highlights key arguments, evidence-based rebuttals, and controversial points raised by both sides. This output improves transparency and interpretability, supporting both machine-side evaluation and human-side engagement. Compared to the D2D, which emphasizes multi-dimensional scoring, ED2D focuses on concise and interpretable decisions. This structure enables downstream applications in user-facing debunking interfaces and facilitates empirical analysis of persuasive effects on belief and behavior.
Misinformation Detection
Experimental Setup
Dataset. We conduct experiments on three datasets: two publicly available benchmarks and one newly constructed resource. The public datasets include Weibo21 (Nan2021) and FakeNewsDataset (perez-rosas-etal-2018-automatic). In addition, we collect Snopes25, a new benchmark compiled from fact-checked real-world claims by professional editors on Snopes, covering the period from January to June 2025. This period is selected to follow the GPT-4o training cutoff to reduce the risk of data leakage from memorized knowledge. Each claim in Snopes25 is annotated as either True or False and paired with an expert-written fact-checking article. Table 1 summarizes the key statistics of the datasets.
Dataset Fake Real Total
Weibo21 2,373 2,461 4,843
FakeNewsDataset 466 466 932
Snopes25 252 196 448
Table 1: Statistics of the three datasets
@@
  • BERT (devlin-etal-2019-bert): A fine-tuned BERT-base model for binary classification.
  • RoBERTa (liu2019robertarobustlyoptimizedbert): A fine-tuned RoBERTa-base model with the same setup as BERT, serving as a stronger discriminative baseline.
@@
  • Chain-of-Thought (CoT) (Weicot): The LLM is prompted to generate a step-by-step reasoning trace before producing a final prediction.
  • Self-Reflect (SR) (madaan2023selfrefine): The model iteratively critiques and revises its own outputs until a convergence criterion is met, typically based on self-evaluation of quality or confidence.
@@
  • Debate-to-Detect (D2D) (han2025d2d): A MAD framework with the same five structured stages. Agents are assigned domain-specific profiles and fixed stances but do not have access to external evidence.
In addition, for each LLM-based baseline, we implement a variant augmented with similar evidence retrieval mechanism. This allows us to assess the contribution of external factual evidence to model performance across different prompting strategies.
@@
Table 2 summarizes the performance of all evaluated methods across the three benchmark datasets. We report standard classification metrics including accuracy (Acc), precision (Prec), recall (Rec), and F1-score (F1). Across all datasets and metrics, ED2D achieves the best performance, significantly outperforming both traditional fine-tuned transformers (BERT and RoBERTa) and recent prompting-based approaches such as CoT, SR, and SMAD. While other deep learning models also yield competitive results, their black-box nature severely limits interpretability. These findings support the conclusion that structured debate, when augmented with external factual evidence, leads to more accurate and reliable misinformation detection. Notably, ED2D consistently outperforms other methods on Snopes25, suggesting strong generalization to real-world claims beyond the model’s training cutoff.
@@
Evidence-based grounding improves performance. For each LLM-based method, we implement a variant that incorporates retrieved external evidence as additional context. Across all datasets and methods, this grounding consistently enhances performance. Simpler prompting methods show the strongest effect, as ZS yields the largest gain and underscores the role of factual grounding in LLM-based reasoning.
From D2D to ED2D. While D2D benefits from structured multi-agent interactions, ED2D introduces explicit evidence retrieval during the free debate stage. This enhancement yields consistent improvements of 2–3 points across all evaluation metrics. Overall, ED2D provides a robust and interpretable architecture for misinformation detection by combining the strengths of multi-agent deliberation and evidence-based reasoning. Its consistent outperformance across diverse benchmarks positions it as a strong candidate for integration into real-world fact-checking and misinformation mitigation workflows.
@@
Figure 1 presents a case from Snopes25 on the claim that pathogens can be released into the air when a toilet is flushed without a closed lid. The affirmative team grounded their argument in scientific summaries retrieved from Wikipedia, highlighting the formation of bioaerosols during flushing and the importance of addressing preventable hygiene risks in public settings. In contrast, the negative team questioned the real-world relevance of these findings, emphasizing the lack of direct causal evidence and arguing against diverting attention from established transmission pathways. The judges ruled in favor of the affirmative, citing stronger evidence synthesis, clearer reasoning, and appropriate use of the precautionary principle. This case demonstrates how structured debate can facilitate interpretable, evidence-based evaluation of contested health claims. The full debate transcript is provided in Appendix B.
To support real-world deployment, we developed a public-facing community website for ED2D, as illustrated in Figure 3. The platform enables users to explore structured debates generated from custom claims. By making the underlying debate process transparent, the website aims to strengthen user resilience to misinformation. A demonstration video is provided in the supplementary materials.
@@
Table 3 reports group-level means for factual accuracy and user ratings of belief, willingness to share, and emotional alignment on a 7-point Likert scale. To assess the reliability of observed effects, we employ mixed-effects regression models with random intercepts for both participants and claims to account for repeated measures. Binary classification accuracy is analyzed using logistic mixed-effects regression, while Likert-scale ratings (belief, willingness to share, emotional alignment) are evaluated using linear mixed-effects regression. Post-hoc pairwise comparisons with Tukey adjustment reveal statistically significant differences across experimental conditions (all $p<0.05$) in both objective accuracy and subjective belief-related measures.
RQ2: Persuasiveness When ED2D is Correct. When ED2D generates correct classifications, its explanations prove notably persuasive and match the effectiveness of expert fact-checking. Compared to participants without additional explanation, those exposed to ED2D or Snopes are more likely to accurately identify the truthfulness of both false and true claims. These explanations also reduce belief in false claims and increase belief in true ones. The joint presentation of ED2D and Snopes yields the strongest effects, suggesting a cumulative benefit when human and AI explanations align. In addition to improving factual accuracy, these explanations reduce participants’ willingness to share misinformation and weaken emotional agreement with false claims, while enhancing alignment and sharing intent for true claims.
@@
RQ3: Misleading Persuasion When ED2D is Incorrect. When ED2D generates incorrect judgments, its explanations distort user beliefs. For false claims misclassified as true, participants who receive the ED2D explanation are more likely to believe and share the misinformation. In contrast, Snopes explanations promote more accurate beliefs and reduce willingness to share. When both explanations are shown together, the misleading influence of ED2D partially offsets the corrective effect of Snopes. A similar pattern emerges for true claims incorrectly labeled as false: ED2D’s explanation reduces belief in accurate information and lowers judgment accuracy, while Snopes helps preserve more accurate perceptions. These results suggest that persuasive but incorrect AI explanations can undermine human fact-checking, even when both are presented side by side.
@@
Figure 4 presents participants’ judgment accuracy under different claim domain and claim timeliness. Under RQ2, ED2D explanations consistently enhance user accuracy across all domains. Compared to the Control condition, ED2D substantially improves performance in categories such as politics, science, and culture, with accuracy levels approaching those achieved by human expert. Similar improvements are observed across both current-event and static knowledge claims, suggesting that ED2D retains persuasive strength regardless of temporal context.
@@
In contrast, under RQ3, when ED2D provides incorrect judgments, its explanations systematically reduce user accuracy across all domains. The decline is particularly pronounced in politicized and entertainment-related content, where users appear more susceptible to the model’s misleading reasoning. Timeliness analysis reveals that current claims are especially vulnerable to misinformation when misclassified by ED2D, whereas non-current claims, such as knowledge-based items, show slightly greater resistance.
@@
To assess whether prior exposure to ED2D explanations facilitates independent misinformation detection, we implement a post-test under the RQ2 condition. After completing the main task, participants from the Control, ED2D, and Snopes groups evaluate a new set of 10 unseen claims (5 true, 5 false) without any accompanying explanations, relying solely on the claim text to make veracity judgments.
Table 4: Post-test accuracy on new claims under RQ2.
@@
As shown in Table 4, all groups exhibit improved post-test accuracy relative to baseline performance in the original Control condition. Notably, participants previously exposed to ED2D explanations demonstrate significant gains, particularly in identifying false claims, suggesting that engagement with evidence-based AI explanations may strengthen transferable reasoning ability. Participants in the Snopes condition achieve the highest post-test accuracy overall, underscoring the enduring advantage of expert-authored fact-checks in fostering epistemic vigilance and promoting effective misinformation detection strategies. These results highlight the potential of MAD debunking to support users’ evaluation of online information.

### v2

@@ Abstract
Multi-agent debate (MAD) frameworks have emerged as promising approaches for misinformation detection by simulating adversarial reasoning. While prior work has focused on detection accuracy, the importance of helping users understand the reasoning behind factual judgments has been overlooked. The debate transcripts generated during MAD offer a rich but underutilized resource for transparent reasoning. In this study, we introduce ED2D, an evidence-based MAD framework that extends previous approach by incorporating factual evidence retrieval. More importantly, ED2D is designed not only as a detection framework but also as a persuasive multi-agent system aimed at correcting user beliefs and discouraging misinformation sharing. We compare the persuasive effects of ED2D-generated debunking transcripts with those authored by human experts. Results demonstrate that ED2D outperforms existing baselines across three misinformation detection benchmarks. When ED2D generates correct predictions, its debunking transcripts exhibit persuasive effects comparable to those of human experts; However, when ED2D misclassifies, its accompanying explanations may inadvertently reinforce users’ misconceptions, even when presented alongside accurate human explanations. Our findings highlight both the promise and the potential risks of deploying MAD systems for misinformation intervention. We further develop a public community website to help users explore ED2D, fostering transparency, critical thinking, and collaborative fact-checking.
@@ Introduction
Misinformation presents a persistent threat to online discourse, public trust, and democratic institutions (Ansari and Alam 2025; Schmitt et al. 2025). In response, researchers have developed a range of computational approaches to automatically detect misleading content. Among these, multi-agent debate (MAD) frameworks have recently gained attention for their ability to simulate adversarial reasoning (Liang et al. 2024), where large language model (LLM) agents engage in structured argumentation to expose factual inconsistencies. These multi-agent systems leverage the complementary strengths of competing perspectives and often yield more robust judgments than a single classifier (Li et al. 2024; Zhang et al. 2025a).
Previous studies employing MAD frameworks for post-verification tasks that focus on identifying the falsehood of claims. However, in the real world, users are active reasoners rather than passive recipients, requiring persuasive explanations and the capacity to resist future misinformation(Danry et al. 2025). Simply labeling a claim as false is insufficient to foster resilience against misinformation (Lyu et al. 2025). Drawing on the adage that the Truth Becomes Clearer Through Debate (Han et al. 2025b; Liu et al. 2025), we argue that effective misinformation interventions should prioritize transparent reasoning, evidence grounding, and persuasive debunking.
In this study, we introduce ED2D, an evidence-based MAD framework that extends the Debate-to-Detect (D2D) framework (Han et al. 2025b) by incorporating an evidence retrieval module. It automatically identifies key entities and concepts within a claim, retrieves factual information from external sources, and assesses the stance of the retrieved evidence to the claim. By integrating verifiable facts into the argumentative process, ED2D mitigates hallucinations and achieves superior performance on three standard misinformation detection datasets.
We further evaluate ED2D’s persuasive effect on human beliefs using Snopes25, a real-world benchmark consisting of fact-checks authored by professional editors from January to June 2025. It is collected after the training cut-off of GPT-4o, thereby reducing the risk of data leakage and better reflecting current misinformation trends. Using Snopes25, we compare the effects of ED2D-generated debunks with human-expert reports on belief correction and sharing intention. We also analyze failure cases in which ED2D misjudges claims yet still successfully persuades users, highlighting potential risks associated with the deployment of MAD systems.
In summary, our study makes the following contributions:
@@
Early approaches focused on fine-tuning pretrained transformers such as BERT for binary veracity classification tasks (Li et al. 2021; Zhu et al. 2025). While these approaches perform well on benchmark datasets, they often lack interpretability and generalize poorly to novel claims (Pelrine et al. 2023; Han and Tang 2026). To overcome these limitations, recent work has shifted toward prompting-based paradigms using LLMs. Zero-shot and few-shot prompting enable flexible adaptation to misinformation detection tasks without the need for explicit fine-tuning (Wu et al. 2023). Moreover, techniques such as Chain-of-Thought (Wei et al. 2022) and Self-Reflection (Madaan et al. 2023) enhance reasoning transparency by enabling models to decompose complex judgments into intermediate steps or engage in iterative self-critique. However, as these approaches depend entirely on internal model knowledge, they are prone to hallucinations and factual errors (Huang et al. 2025; Yu et al. 2025). To address these reliability concerns, an emerging line of research incorporates external evidence into the reasoning process. For example, Retrieval-Augmented Generation (RAG) has been widely adopted to ground LLM outputs in verifiable content. Nevertheless, most existing RAG systems still rely on a fixed knowledge base, which limits their effectiveness in open-domain misinformation detection scenarios(Wang et al. 2025).
@@
MAD frameworks have gained traction as an approach to improve the deliberative quality of LLM outputs by simulating adversarial interactions between agents with opposing views (Du et al. 2024; Liang et al. 2024; Li et al. 2024). Drawing inspiration from human deliberation and dialectical reasoning, such frameworks allow agents to challenge assumptions, correct errors, and elicit deeper justifications(Zhang et al. 2025b).
@@
Beyond accurate classification, effective misinformation mitigation requires persuasive interventions that influence user beliefs and behaviors (Ecker et al. 2022). Research in cognitive psychology has explored the design of corrective messages, demonstrating that informative and clear explanations can reduce belief in falsehoods and discourage the sharing intention (Jahanbakhsh et al. 2021). Factors such as emotion, epistemic trust, and narrative coherence have been shown to moderate the effectiveness of fact-checking efforts (Pennycook et al. 2021; Scherer et al. 2021).
Professional fact-checking organizations, such as Snopes and PolitiFact, consistently produce structured and evidence-based rebuttals that represent the current gold standard in combating online misinformation. However, expert-written fact-checks are often costly and inherently lack scalability. Consequently, there is growing interest in automating persuasive debunking using LLMs (Salvi et al. 2025; Schoenegger et al. 2025). However, few studies have directly compared the persuasive efficacy of AI-generated versus expert-written explanations.
@@
Figure 1 presents the architecture of ED2D. The framework builds upon the five-stage debate structure: Opening Statement, Rebuttal, Free Debate, Closing Statement, and Judgment. At the core of the system lies the Agent Layer, comprising two debating teams—the Affirmative and the Negative—each consisting of four agents. These agents are assigned domain-specific profiles relevant to the input and are fixed to either a “True” or “Fake” stance. Within each team, agents collaborate to construct coherent arguments that support or refute the veracity of the claim.
The debate is evaluated by a panel of judge agents, who observe the full dialogue and score it across five dimensions: Factuality, Source Reliability, Reasoning Quality, Clarity, and Ethical Considerations.Each dimension is evaluated using a complementary scoring scheme in which paired scores sum to seven, thereby precluding any possibility of a tie. The aggregated scores result in a definitive classification of the claim as either REAL or FAKE.
The Orchestrator Layer governs the overall debate process, assigning roles, scheduling turns across the five structured stages, and maintaining a compressed shared dialogue memory. To mitigate the context-length limitations of LLMs, the system performs stage-wise context compression, distilling salient information into concise summaries that guide subsequent reasoning stages. This mechanism ensures continuity and coherence throughout the multi-turn debate.
@@
Evidence Retrieval and Integration
  1. 1.
    Entity and Relation Extraction: Using in-context prompting with LLMs, the system identifies up to five salient entities or concepts from the input claim.
  2. 2.
  3. 3.
  4. 4.
Judgment and Output
In the final stage, ED2D generates two outputs: a binary veracity label and a structured summary of the debate. The summary highlights key arguments, evidence-based rebuttals, and controversial points raised by both sides. The output enhances transparency and interpretability, supporting both machine-side evaluation and human-side engagement. In contrast to D2D, which emphasizes multi-dimensional scoring, ED2D prioritizes concise and interpretable decision-making. Such a design enables downstream integration into user-facing debunking interfaces and facilitates empirical analyses of persuasive effects on users’ beliefs and behaviors.
Misinformation Detection
Experimental Setup
Dataset. We conduct experiments on three datasets: two publicly available benchmarks and one newly constructed resource. The public datasets include Weibo21 (Nan et al. 2021) and FakeNewsDataset (Pérez-Rosas et al. 2018). In addition, we collect Snopes25, a new benchmark compiled from fact-checked real-world claims by professional editors on Snopes, covering the period from January to June 2025. This period is selected to follow the GPT-4o training cutoff to reduce the risk of data leakage from memorized knowledge. Each claim in Snopes25 is annotated as either True or False and paired with an expert-written fact-checking article. Table 2 summarizes the key statistics of the datasets.
Dataset Fake Real Total
Weibo21 2,373 2,461 4,843
FakeNewsDataset 466 466 932
Snopes25 252 196 448
Table 2: Statistics of the three datasets
@@
  • BERT (Devlin et al. 2019): A fine-tuned BERT-base model for binary classification.
  • RoBERTa (Liu et al. 2019): A fine-tuned RoBERTa-base model with the same setup as BERT, serving as a stronger discriminative baseline.
@@
  • Chain-of-Thought (CoT) (Wei et al. 2022): The LLM is prompted to generate a step-by-step reasoning trace before producing a final prediction.
  • Self-Reflect (SR) (Madaan et al. 2023): The model iteratively critiques and revises its own outputs until a convergence criterion is met, typically based on self-evaluation of quality or confidence.
@@
  • Debate-to-Detect (D2D) (Han et al. 2025b): A MAD framework with the same five structured stages. Agents are assigned domain-specific profiles and fixed stances but do not have access to external evidence.
In addition, each LLM-based baseline is further extended with a comparable evidence retrieval module, enabling evaluation of the impact of external factual evidence across prompting strategies.
@@
Table 1 summarizes model performance across the three benchmark datasets using standard metrics—accuracy (Acc), precision (Prec), recall (Rec), and F1-score (F1). ED2D achieves the strongest results on every dataset and metric, substantially outperforming both fine-tuned transformers (BERT, RoBERTa) and prompting-based methods such as CoT, SR, and SMAD. Although several deep learning baselines yield competitive accuracy, their limited interpretability constrains practical deployment. These results demonstrate that structured debate, when augmented with external factual evidence, produces more accurate and reliable misinformation detection. ED2D’s consistent superiority on Snopes25 further indicates strong generalization to real-world, post-training claims.
@@
Evidence-based grounding improves performance. For each LLM-based method, we evaluate a variant that incorporates retrieved evidence as contextual input. Grounding yields uniform improvements across all methods, with the largest gains observed for simpler prompting approaches such as ZS, highlighting the central role of factual support in LLM reasoning.
From D2D to ED2D. While D2D benefits from structured multi-agent deliberation, ED2D’s integration of explicit evidence retrieval during the free-debate phase produces consistent 2–3 point improvements across all metrics. By combining multi-agent argumentation with evidence-based reasoning, ED2D offers a robust and interpretable architecture well suited to real-world fact-checking and misinformation mitigation workflows.
@@
Figure 2 illustrates a representative Snopes25 case concerning whether flushing a toilet with the lid open releases pathogens into the air. The affirmative team anchored its argument in scientific summaries retrieved from Wikipedia, emphasizing the formation of bioaerosols and the relevance of preventative hygiene measures. The negative team challenged the practical significance of these findings, citing limited causal evidence and cautioning against distraction from more established transmission routes. The judges favored the affirmative team on the basis of stronger evidence synthesis, clearer reasoning, and appropriate reliance on the precautionary principle.
To support real-world use, we additionally developed a publicly accessible ED2D community platform that allows users to generate and inspect structured debates for custom claims, as illustrated in Figure 3. By exposing the full deliberation process, the system aims to improve user resilience to misinformation. A demonstration video is included in the supplementary materials.
@@
Table 3 presents group-level means for factual accuracy and user ratings of belief, willingness to share, and emotional alignment on a 7-point Likert scale. We assess effect reliability using mixed-effects regression with random intercepts for participants and claims to account for repeated observations. Binary accuracy is modeled with logistic mixed-effects regression, and Likert-scale outcomes with linear mixed-effects regression. Tukey-adjusted post-hoc comparisons indicate significant differences across conditions (all $p<0.05$) in both objective accuracy and subjective belief-related measures.
RQ2: Persuasiveness When ED2D Is Correct. When ED2D produces correct labels, its explanations are strongly persuasive and comparable to expert fact-checking. Relative to the no-explanation baseline, ED2D and Snopes both improve participants’ truthfulness judgments for true and false claims, lower belief in misinformation, and strengthen belief in accurate content. These explanations also reduce willingness to share false claims and weaken emotional alignment with them, while increasing alignment and sharing intent for true claims.
@@
RQ3: Misleading Persuasion When ED2D Is Incorrect. When ED2D provides incorrect labels, its explanations systematically distort user judgment. For false claims misclassified as true, ED2D increases belief and sharing intention, whereas Snopes mitigates belief in misinformation and reduces sharing. When both explanations are shown, ED2D’s misleading influence partially counteracts Snopes’ corrective effect. For true claims misclassified as false, ED2D suppresses belief in accurate information and lowers judgment accuracy, while Snopes helps preserve more appropriate evaluations. Overall, these findings show that persuasive but erroneous AI explanations can undermine human fact-checking—even when presented alongside authoritative guidance.
@@
Figure 4 reports participants’ judgment accuracy across claim domains and timeliness conditions. Under RQ2, ED2D explanations consistently increase accuracy in every domain. Relative to the Control group, ED2D yields marked gains in politics, science, and culture, achieving accuracy levels close to those of human experts. Comparable improvements appear for both current-event and static knowledge claims, indicating that ED2D’s persuasive effectiveness is largely insensitive to temporal context.
@@
Under RQ3, when ED2D produces incorrect labels, its explanations uniformly depress accuracy across domains. The reduction is most pronounced for politicized and entertainment-related content, where participants are more easily influenced by misleading reasoning. Timeliness analysis further shows that current-event claims are especially susceptible to misclassification-driven misinformation effects, whereas static, knowledge-based claims exhibit slightly greater robustness.
@@
To evaluate whether exposure to ED2D explanations improves participants’ ability to independently detect misinformation, we conduct a post-test within the RQ2 condition. After completing the main task, participants from the Control, ED2D, and Snopes groups assess a new set of ten claims (five true and five false) without any explanatory material, basing their veracity judgments solely on the claim text.
@@
Table 4: Post-test accuracy on new claims under RQ2.
As shown in Table 4, all groups exhibit higher post-test accuracy compared with baseline performance in the original Control condition. Participants previously exposed to ED2D explanations show substantial improvements, especially in detecting false claims, indicating that engagement with evidence-based AI explanations can enhance transferable reasoning skills. Participants in the Snopes condition attain the highest overall accuracy, reaffirming the lasting benefit of expert fact-checks in promoting epistemic vigilance and effective misinformation detection. Taken together, these findings suggest that MAD-style debunking can strengthen users’ ability to evaluate online information even in the absence of explicit explanatory support.